Good evening! Hope you have a grate day!
Please, help me. How these logarithms could be solved?

1) 2. 952 = 6* loge 2.952=nloge 16 n= 2 4) 2. /6P - m lope 1.168 = m loge 2 168 = mloge?
Answer #1

2 - Qs 2 - 1 де (2 - 952) = l»). ! ..3 - Ljе (2 - 252) = м Lo3 4-°bly - 1 - 09,2. S = 72-7726 - и - 1- oks 2. + 6 - б. 34о4